OpenStudy (anonymous):

Two riders are in a bike race. The second rider starts half an hour late but rides 4 miles per hour faster than the first rider. If they both finish 3 hours after the race starts, how fast are they each riding?

OpenStudy (anonymous):

Let's call x the amount of time the first rider has been on the road. Let's also call v the velocity of the first rider. An equation for the distance for the first rider has traveled is \[D_1=vx\] The second rider has an initial start time of (x-0.5) because he starts half an hour later than the first rider. The second rider's velocity is 4 mph faster than the first rider, or (v+4). This gives us an equation for the distance the second rider has traveled \[D_2=(v+4)(x-0.5)\] They finish 3 hours after the race starts (x=3), which also means they traveled the same distance, so \[D_1=D_2\]\[vx=(v+4)(x-0.5)\]\[v(3)=(v+4)(3-0.5)\]\[3v=(v+4)2.5\]Distribute \[3v=2.5v+10\]Subtract 2.5v from both sides to get v on one side. \[0.5v=10\]Multiply by 2 to get 1v \[v=20\] The second rider is riding 4 mph faster, so the second rider is riding 24 mph.

OpenStudy (rosey):

i can't help but u should b a fan of lil ol me :)

OpenStudy (anonymous):

Who should?